OpenStudy (supersquidsisters):

help i cant figure this out? Two gears are connected and are rotating simultaneously. The smaller gear has a radius of 4 inches, and the larger gear has a radius of 7 inches. Part 1: What is the angle measure, in degrees and rounded to the nearest tenth, through which the larger gear has rotated when the smaller gear has made one complete rotation? Part 2: How many rotations will the smaller gear make during one complete rotation of the larger gear?

OpenStudy (supersquidsisters):

heres the picture of the gears

OpenStudy (radar):

Calculate how far a point would travel around the perimeter (circumference) of the small gear for one revolution. In other words calculate the perimeter of the smaller gear.

OpenStudy (jadeishere):

The first step in finding the answer is to find the circumference of each gear. \[circumference = 2PIr\] 2(3)pi = 6pi 2(7)pi = 14pi

OpenStudy (jadeishere):

(I couldn't find the pi symbol...)

OpenStudy (jadeishere):

@radar Do you want to take this question?

OpenStudy (radar):

Use 3.5 instead of 3. No you have the student on the right path.

OpenStudy (jadeishere):

Oh, I thought it was 3 not 4. Oops xD Sorry!! I read that wrong

OpenStudy (supersquidsisters):

@Jadeishere so 2(3.5)pi = 6pi like @radar said?

OpenStudy (radar):

I may have goofed, use the diameter or two radius.

OpenStudy (jadeishere):

Well, let's change it to 4, cause I read it as 3. My mind does crazy things, I apologize. Kay, backtracking. \[2\pi(4) = 8\pi\] \[2\pi(7) = 14\pi\]

OpenStudy (jadeishere):

So then, to get the angle measure in degrees, we have to divide the smaller gears circumference by the larger gears circumference then multiply it by 360 degrees. \[\frac{ 8\pi }{ 14\pi } * 360\]

OpenStudy (jadeishere):

That can be further simplified to \[(\frac{ 8 }{ 14 })*360\]

OpenStudy (jadeishere):

You can actually just simplify that right back down to \[(\frac{ 4 }{ 7 }) * 360\] That may seem like a bit of a circle, but make sure to take every necessary step.

OpenStudy (jadeishere):

The number you get for that, will be the answer, in degrees, to Part 1.

OpenStudy (supersquidsisters):

@Jadeishere ok so i got 205.7142857 when i solved the equation

OpenStudy (jadeishere):

Then, like it said, round to the nearest tenth. :)

OpenStudy (supersquidsisters):

@Jadeishere oh so 206?

OpenStudy (radar):

@adeishere Love your profile.

OpenStudy (jadeishere):

Tenth, not whole number :) A tenth is any decimal from 0.1 - 0.9 So it'd still be 205.7_____ Is 0.7 the nearest tenth, or can you round up from there?

OpenStudy (jadeishere):

Thanks @radar

OpenStudy (supersquidsisters):

@Jadeishere ??? um i guess you can ??????????

OpenStudy (jadeishere):

Rounding up or rounding down; Up = 0.75+ down = 0.74- So is the number directly to the right of 205.7 greater than 5 or less than 5?

OpenStudy (supersquidsisters):

@Jadeishere oh it was less than

OpenStudy (jadeishere):

So we round down to 205.7 :) The answer to Part 1 is 205.7 degrees :D Now for Part 2. This one is quite a bit simpler than than the previous Part. This time, all you have to do is divide the larger radius by the smaller radius :)

OpenStudy (supersquidsisters):

@Jadeishere ah ok!! gimmie a bit d:

OpenStudy (supersquidsisters):

@Jadeishere 1.75?

OpenStudy (jadeishere):

Yes! Great job! :D

OpenStudy (jadeishere):

Great work @supersquidsisters I'm glad I was able to help you! :)

OpenStudy (supersquidsisters):

@Jadeishere thank you soo much!!! you were such great help!!

OpenStudy (jadeishere):

I'm glad to hear :)